OpenStudy (anonymous):

if the current in a wire is at 0.5 A and the voltage is measured at 100 V what is the resistance in the wire

OpenStudy (anonymous):

E=IR 100 = 0.5 * R R=100/0.5 = 200

OpenStudy (anonymous):

heywood is correct!:D Great job and Good luck:D